OpenStudy (anonymous):

Find the indefinite integral (3x-7)^-6dx

hartnn (hartnn):

you can substitute u = 3x-7 then, du=... ?

OpenStudy (anonymous):

right, it would be ∫(3x-7) ^ 6 dx , substitute u = 3x -7 , du = 3dx , du/3 = dx . 1/3 *∫u^-6 , 1/3* -u/5 ^ -5 + C .

OpenStudy (anonymous):

sorry , 1/3 * (u)^-5/-5 + c , then replace u.

OpenStudy (anonymous):

Thank you! Trying to wrap my head around it!

OpenStudy (anonymous):

medal -.- ..